OpenStudy (freakzter):

Match item in Column A to an answer in Column B. What is the distance between the two points on a number line? Column A 1. X=4. Y= -17 2. P=23, Q=8 3. U=-1. V=-14 4. J=-3, K=18 Column B A. 13 B. 15 C. 21

jimthompson5910 (jim_thompson5910):

Draw a number line. Plot the points 4 and -7 on it. How many spaces are between the two numbers?

OpenStudy (freakzter):

21 (I messed up it isn't 7, it's 17)

jimthompson5910 (jim_thompson5910):

oh so X = 4 and Y = -17 ?

OpenStudy (freakzter):

Yes, sorry

jimthompson5910 (jim_thompson5910):

yes, the distance between the two values is 21 units

OpenStudy (freakzter):

How about 2. and 4.?

jimthompson5910 (jim_thompson5910):

in #2, the two numbers are positive, so you can subtract 23 - 8 = 15

OpenStudy (freakzter):

Oh, makes sense thanks

jimthompson5910 (jim_thompson5910):

you can also subtract in #1 x - y = 4 - (-17) = 4 + 17 = 21

OpenStudy (freakzter):

Oh thanks

jimthompson5910 (jim_thompson5910):

no problem